IPhoneDude120:

4(x-12)+19-19=-89-19 4(x-12)=-108 4(x-12)/4 = -108/4 x-12=-27 x-12+12=-27+12 x=-?

IPhoneDude120:

no its not! hahhah 4(x-12)+19-19=-89-19 4(x-12)=-108 4(x-12)/4 = -108/4 x-12=-27 x-12+12=-27+12 x=-15